OpenStudy (batman19991):

The least common multiple of 20, 24, and 45 is _____. 30 180 360 21,600

OpenStudy (solomonzelman):

20=4×5 24=4×3×2 45=5×3×3 so you need to use the 2 once, the 3 twice, 4 once, and 5 once.

OpenStudy (batman19991):

what

OpenStudy (batman19991):

@misssunshinexxoxo

OpenStudy (australopithecus):

https://www.mathsisfun.com/least-common-multiple.html this explains the concept of least common multiples pretty well

OpenStudy (batman19991):

ok thanks

OpenStudy (anonymous):